OpenStudy (anonymous):

Anyone?: \[\int\limits12t^{7}-t ^{2}+3dt \]

OpenStudy (anonymous):

come on, lets be seriouse

OpenStudy (saifoo.khan):

\[\frac{12}{8}t^8 - \frac{t^3}{3}+3x\]

OpenStudy (saifoo.khan):

\[\frac{3}{2}t^8 - \frac{t^3}{3}+3x\]

OpenStudy (anonymous):

you forgot something^

OpenStudy (mimi_x3):

um, i thnk its "t" and you for + C

OpenStudy (mimi_x3):

forgot*

OpenStudy (mertsj):

+c

OpenStudy (saifoo.khan):

And +C

OpenStudy (anonymous):

yep... that's what I got... Thanks guys :)

OpenStudy (saifoo.khan):

welcome.